Ammonium nitrate (NH₄NO₃), an important fertilizer, produces dinitrogen oxide (N₂O) gas and H₂O when it decomposes. Determine the mass of H₂O produced from the decomposition of 25.0 grams of solid NH₄NO₃. Reaction: NH₄NO₃ -> N₂O + 2 H₂O



1
Expert's answer
2021-10-26T05:34:12-0400

M(NH4NO3) = 80.04 g/mol

m(NH4NO3) "= \frac{25.0}{80.04} = 0.312 \;mol"

According to the reaction:

n(H2O) = 2n(NH4NO3) "= 2 \times 0.312 = 0.624 \;mol"

M(H2O) = 18.01 g/mol

m(H2O) "= 0.624 \times 18.01 = 11.25 \;g"

Answer: 11.25 g